Grade 12Algebra

The number 2, a, b are in arithmetic progression and a, b, 9 are in geometric progression, How can I find the values of a and b?

Junior Huriavi
8 Years agoGrade 12

2 Answers

Swaraj
8 Years ago
a=2 + d
b=2 + 2d
b=a x r
9=b x r
 
4 variables 4 equations . solving them you will get a=4  and b=6
And also d=2 and r=1.5
krishna bhagat
8 Years ago
 
Not so tough using equations
a=2 + d
b=2 + 2d
b=a x r
9=b x r
a=4
b=6
d=2
r=1.5
 
 a=4  and b=6
And also d=2 and r=1.5